Flavonoids are plant-based phytochemicals with cardiovascular protective properties. Common sources of flavonoids are berries, tree fruits, some nuts, and beans. The incidence of cardiovascular disease hospitalizations was compared between adult men age 35-50 years who consumed flavonoids in their diet compared to those who do not consume flavonoids. The outcome measure was hospital discharges for cardiovascular diseases. See the table below for the data collected.

Eats Flavonoids

Does Not Eat Flavonoids

Total

Number of Hospitalizations

50

90

140

Person Months

900

1200

2100

  1. Describe the appropriate 2 x 2 table of the data from this study. Include the column and row names and the numbers that would appear in each table cell.
Answer #1

By the table given above the following parameters has to be calculated :-

True positive - 50

True negative - 1200

False positive -90

False negative - 900

Specificity ratio = TP/(TP+ FN)

= 50/(50+900)

= 50/950

= 0.05%

Sensitivity ratio= TN/(TN+ FP)

= 1200/(1200+90)

= 1200/1290

= 0.93%

Positive predictive value = TP/(TP+FP)

= 50/(50+90)

= 50/140

= 0.3%

Negative predictive value = TN/(TN+FN)

= 1200/(1200+900)

= 1200/2100

= 0.5%